5 step process
Less Tar, more Nicotine? Many cigarette manufacturers market their cigarettes by saying they have low tar. Some people believe that in order to compensate for the low tar and still keep sales high, the manufacturers are increasing the average amount of nicotine in their cigarettes. A group of quality assurance personnel from the FDA sampled 41 low tar cigarettes from brand XYZ. The sample produced an average of 1.5 mg of nicotine with s = 0.33 mg. The national allowable average of nicotine in low tar cigarettes is 1.06 mg. Test the claim that the XYZ brand has a mean amount of nicotine that is greater than the allowable limit. (use a = 0.05)

Answer by stanbon(75887)

A group of quality assurance personnel from the FDA sampled 41 low tar cigarettes from brand XYZ. The sample produced an average of 1.5 mg of nicotine with s = 0.33 mg. The national allowable average of nicotine in low tar cigarettes is 1.06 mg. Test the claim that the XYZ brand has a mean amount of nicotine that is greater than the allowable limit. (use a = 0.05)
--------------
Ho: u = 1.06
Ha: u > 1.06
---
x-bar = 1.5
---
t(1.5) = (1.5-1.06)/[0.33/sqrt(41)] = 8.5375
---
p-value = P(t> 8.5375 when df=40) =7.5x10^-11
----
Conclusion: Since the p-value is less than 5%,
reject Ho.
================
Cheers,
Stan H.